# Native Design Bias: Studying the Impact of English Nativeness on Language Model Performance

### v2

@@ Appendix I Results Sonnet different languages
Appendix K Prompt guess native
For this prompt, the number of identified correct predictions by the models was gathered automatically. Table 16 shows an overview of the correctly identified native and non native prompts for the entire dataset. In total there are 4264 native prompts and 8255 non-native prompts in our dataset. Many of the incorrectly predicted prompts, where in fact indecisive guesses or when the model was in doubt.
amount guess native correctly amount guess non-native correctly
GPT3.5 350 3080
GPT4o 114 436
Haiku 164 914
Sonnet 87 304
qwen 7B 95 152
Table 16: Amount of times the user’s nativeness was correctly guessed by the different models.
Appendix L Generative results
Tables 17 and 18 show the full performance metrics for the different experiments throughout the paper averaged over three runs including the standard deviation. In bold, we indicated the significant differences where (p<0.05). We tested this using a Wilcoxon signed-rank test over all runs and metrics.
GPT3.5 GPT4o
native non-native native non-native
BLEU $0.44\pm 0.00$ $0.44\pm 0.00$ $0.47\pm 0.00$ $0.47\pm 0.00$
ROUGE-1 $0.39\pm 0.01$ $0.39\pm 0.00$ $0.40\pm 0.00$ $0.41\pm 0.00$
ROUGE-2 $0.15\pm 0.01$ $0.15\pm 0.01$ $0.16\pm 0.01$ $0.16\pm 0.00$
ROUGE-L $0.28\pm 0.01$ $0.29\pm 0.00$ $0.29\pm 0.00$ $0.29\pm 0.00$
BERTscore precision $0.72\pm 0.00$ $0.72\pm 0.00$ $0.72\pm 0.00$ $0.72\pm 0.00$
BERTscore recall $0.72\pm 0.00$ $0.72\pm 0.00$ $0.72\pm 0.00$ $0.72\pm 0.00$
BERTscore F1 $0.72\pm 0.00$ $0.72\pm 0.00$ $0.72\pm 0.00$ $0.72\pm 0.00$
BARTscore faithful $-2.79\pm 0.01$ $-2.79\pm 0.01$ $-2.96\pm 0.01$ $-2.97\pm 0.00$
BARTscore precision $-2.73\pm 0.00$ $-2.72\pm 0.01$ $-2.75\pm 0.01$ $-2.75\pm 0.01$
BARTscore recall $-2.73\pm 0.00$ $-2.71\pm 0.01$ $-2.77\pm 0.01$ $-2.76\pm 0.00$
BARTscore F1 $-2.70\pm 0.01$ $-2.69\pm 0.00$ $-2.74\pm 0.01$ $-2.73\pm 0.00$
Haiku Sonnet Qwen 7B
native non-native native non-native native non-native
BLEU $0.45\pm 0.00$ $0.46\pm 0.01$ $0.45\pm 0.00$ $0.45\pm 0.00$ $0.41\pm 0.00$ $0.42\pm 0.01$
ROUGE-1 $0.41\pm 0.01$ $0.42\pm 0.01$ $0.38\pm 0.00$ $0.38\pm 0.01$ $0.37\pm 0.01$ $0.38\pm 0.00$
ROUGE-2 $0.16\pm 0.00$ $0.17\pm 0.01$ $0.13\pm 0.01$ $0.14\pm 0.00$ $0.14\pm 0.01$ $0.14\pm 0.00$
ROUGE-L $0.30\pm 0.01$ $0.30\pm 0.00$ $0.27\pm 0.00$ $0.27\pm 0.00$ $0.26\pm 0.01$ $0.27\pm 0.00$
BERTscore precision $0.72\pm 0.00$ $0.72\pm 0.00$ $0.71\pm 0.00$ $0.71\pm 0.01$ $0.70\pm 0.00$ $0.70\pm 0.00$
BERTscore recall $0.71\pm 0.01$ $0.72\pm 0.00$ $0.71\pm 0.00$ $0.71\pm 0.00$ $0.69\pm 0.01$ $0.70\pm 0.00$
BERTscore F1 $0.72\pm 0.01$ $0.72\pm 0.00$ $0.71\pm 0.00$ $0.71\pm 0.00$ $0.70\pm 0.01$ $0.70\pm 0.00$
BARTscore faithful $-2.89\pm 0.01$ $-2.90\pm 0.01$ $-3.13\pm 0.01$ $-3.12\pm 0.01$ $-3.11\pm 0.02$ $-3.10\pm 0.00$
BARTscore precision $-2.69\pm 0.01$ $-2.70\pm 0.01$ $-2.87\pm 0.01$ $-2.86\pm 0.01$ $-2.9\pm 0.01$ $-2.91\pm 0.01$
BARTscore recall $-2.69\pm 0.01$ $-2.67\pm 0.01$ $-2.88\pm 0.00$ $-2.87\pm 0.00$ $-2.95\pm 0.02$ $-2.93\pm 0.01$
BARTscore F1 $-2.67\pm 0.01$ $-2.66\pm 0.00$ $-2.85\pm 0.01$ $-2.84\pm 0.01$ $-2.9\pm 0.01$ $-2.89\pm 0.01$
Table 17: Overall average generative performance metrics and standard deviation over three runs for all generative tasks per model and group .
GPT3.5 GPT4o
western native not western native western native not western native
BLEU $0.44\pm 0.00$ $0.44\pm 0.00$ $0.47\pm 0.00$ $0.47\pm 0.00$
ROUGE-1 $0.39\pm 0.01$ $0.39\pm 0.00$ $0.40\pm 0.00$ $0.41\pm 0.00$
ROUGE-2 $0.14\pm 0.01$ $0.15\pm 0.00$ $0.16\pm 0.00$ $0.16\pm 0.01$
ROUGE-L $0.28\pm 0.01$ $0.29\pm 0.00$ $0.29\pm 0.00$ $0.29\pm 0.00$
BERTscore precision $0.72\pm 0.00$ $0.72\pm 0.00$ $0.72\pm 0.00$ $0.72\pm 0.00$
BERTscore recall $0.72\pm 0.00$ $0.72\pm 0.00$ $0.72\pm 0.00$ $0.72\pm 0.00$
BERTscore F1 $0.72\pm 0.00$ $0.72\pm 0.00$ $0.72\pm 0.00$ $0.72\pm 0.00$
BARTscore faithful $-2.79\pm 0.01$ $-2.79\pm 0.02$ $-2.97\pm 0.01$ $-2.96\pm 0.01$
BARTscore precision $-2.73\pm 0.01$ $-2.72\pm 0.01$ $-2.74\pm 0.01$ $-2.75\pm 0.00$
BARTscore recall $-2.73\pm 0.00$ $-2.71\pm 0.00$ $-2.76\pm 0.01$ $-2.76\pm 0.00$
BARTscore F1 $-2.71\pm 0.01$ $-2.69\pm 0.00$ $-2.73\pm 0.00$ $-2.73\pm 0.00$
Haiku Sonnet Qwen 7B
western native not western native western native not western native western native not western native
BLEU $0.45\pm 0.00$ $0.46\pm 0.01$ $0.45\pm 0.00$ $0.45\pm 0.00$ $0.41\pm 0.00$ $0.42\pm 0.01$
ROUGE-1 $0.40\pm 0.01$ $0.42\pm 0.01$ $0.38\pm 0.00$ $0.38\pm 0.01$ $0.37\pm 0.01$ $0.38\pm 0.00$
ROUGE-2 $0.16\pm 0.00$ $0.17\pm 0.01$ $0.14\pm 0.01$ $0.14\pm 0.00$ $0.14\pm 0.01$ $0.14\pm 0.00$
ROUGE-L $0.29\pm 0.01$ $0.30\pm 0.00$ $0.27\pm 0.00$ $0.27\pm 0.00$ $0.26\pm 0.01$ $0.27\pm 0.00$
BERTscore precision $0.72\pm 0.00$ $0.72\pm 0.00$ $0.71\pm 0.00$ $0.71\pm 0.01$ $0.70\pm 0.00$ $0.70\pm 0.00$
BERTscore recall $0.71\pm 0.01$ $0.72\pm 0.00$ $0.71\pm 0.00$ $0.71\pm 0.00$ $0.70\pm 0.01$ $0.70\pm 0.00$
BERTscore F1 $0.72\pm 0.01$ $0.72\pm 0.00$ $0.71\pm 0.00$ $0.71\pm 0.00$ $0.70\pm 0.01$ $0.70\pm 0.00$
BARTscore faithful $-2.90\pm 0.01$ $-2.89\pm 0.01$ $-3.13\pm 0.01$ $-3.12\pm 0.01$ $-3.11\pm 0.02$ $-3.10\pm 0.00$
BARTscore precision $-2.70\pm 0.01$ $-2.69\pm 0.01$ $-2.86\pm 0.01$ $-2.86\pm 0.01$ $-2.91\pm 0.01$ $-2.90\pm 0.01$
BARTscore recall $-2.70\pm 0.01$ $-2.67\pm 0.01$ $-2.88\pm 0.00$ $-2.87\pm 0.00$ $-2.96\pm 0.02$ $-2.93\pm 0.01$
BARTscore F1 $-2.67\pm 0.01$ $-2.66\pm 0.00$ $-2.84\pm 0.01$ $-2.84\pm 0.01$ $-2.90\pm 0.01$ $-2.88\pm 0.01$
Table 18: Overall average generative performance metrics and standard deviation over three runs for all generative tasks per model and group.
Table 19 shows the overall average BLEU scores and standard deviation over the three runs per dataset and model. All models perform significantly better the non-native group for CODA19, which includes a medical research article generation task, and all models except Haiku for the Paraphrase dataset. For these two tasks, also all models prefer the non western group, except Sonnet on CODA19. Interestingly, for the StoryCloze, we see that GPT3.5 and GPT4o perform significantly better for both the native and western native groups. For the CNN dailymail dataset, almost no significant differences are found.
Story Cloze CNN dailymail CODA19 Paraphrasing
native non-native native non-native native non-native native non-native
GPT3.5 $0.39\pm 0.01$ $0.38\pm 0.01$ $0.47\pm 0.01$ $0.45\pm 0.01$ $0.55\pm 0.00$ $0.56\pm 0.00$ $0.35\pm 0.01$ $0.36\pm 0.00$
GPT4o $0.42\pm 0.01$ $0.41\pm 0$ $0.53\pm 0.00$ $0.53\pm 0.00$ $0.57\pm 0.00$ $0.58\pm 0.00$ $0.35\pm 0.00$ $0.37\pm 0.00$
Haiku $0.40\pm 0.02$ $0.39\pm 0.03$ $0.50\pm 0.01$ $0.51\pm 0.02$ $0.57\pm 0.00$ $0.58\pm 0.00$ $0.34\pm 0.02$ $0.35\pm 0.03$
Sonnet $0.39\pm 0.00$ $0.39\pm 0.01$ $0.51\pm 0.01$ $0.50\pm 0.00$ $0.57\pm 0.00$ $0.58\pm 0.00$ $0.32\pm 0.00$ $0.33\pm 0.00$
Qwen 7B $0.38\pm 0.01$ $0.38\pm 0.01$ $0.42\pm 0.01$ $0.43\pm 0.01$ $0.52\pm 0.01$ $0.53\pm 0.00$ $0.31\pm 0.01$ $0.32\pm 0.01$
Story Cloze CNN dailymail CODA19 Paraphrasing
Western non Western Western non Western Western non Western Western non Western
GPT3.5 $0.39\pm 0.01$ $0.38\pm 0.00$ $0.47\pm 0.01$ $0.46\pm 0.00$ $0.55\pm 0.00$ $0.56\pm 0.00$ $0.35\pm 0.01$ $0.36\pm 0.00$
GPT4o $0.43\pm 0.00$ $0.41\pm 0$ $0.53\pm 0.01$ $0.53\pm 0.00$ $0.58\pm 0.00$ $0.58\pm 0.00$ $0.35\pm 0.00$ $0.37\pm 0.00$
Haiku $0.39\pm 0.03$ $0.40\pm 0.02$ $0.50\pm 0.01$ $0.51\pm 0.02$ $0.57\pm 0.01$ $0.58\pm 0.00$ $0.33\pm 0.02$ $0.35\pm 0.03$
Sonnet $0.40\pm 0.01$ $0.39\pm 0.00$ $0.51\pm 0.01$ $0.50\pm 0.00$ $0.57\pm 0.01$ $0.58\pm 0.00$ $0.32\pm 0.00$ $0.33\pm 0.00$
Qwen 7B $0.38\pm 0.01$ $0.38\pm 0.01$ $0.42\pm 0.01$ $0.43\pm 0.01$ $0.52\pm 0.00$ $0.53\pm 0.01$ $0.30\pm 0.01$ $0.33\pm 0.00$
Table 19: Average BLEU scores and standard deviation per dataset and model for the native-non-native group and the Western non Western group over the three runs. The statistical analysis is done based on all different performance scores using a Wilcoxon signed-rank test with p-value 0.05. Significant differences are indicated in bold
Appendix M Overall results per prompt
Tables 20, 21, 22, 23, and 24 show an overview of the overall average results per group, model, and prompt.
(non-)native system prompt
correct information wrong information
native non-native native non-native
BLEU $0.42\pm 0.01$ $0.40\pm 0.01$ $0.40\pm 0.00$ $0.42\pm 0.01$
ROUGE-1 $0.38\pm 0.00$ $0.37\pm 0.00$ $0.37\pm 0.00$ $0.38\pm 0.01$
ROUGE-2 $0.14\pm 0.00$ $0.13\pm 0.00$ $0.13\pm 0.00$ $0.14\pm 0.00$
ROUGE-L $0.28\pm 0.00$ $0.28\pm 0.00$ $0.27\pm 0.00$ $0.28\pm 0.00$
BERTscore precision $0.72\pm 0.00$ $0.73\pm 0.00$ $0.72\pm 0.00$ $0.72\pm 0.00$
BERTscore recall $0.71\pm 0.00$ $0.70\pm 0.00$ $0.70\pm 0.00$ $0.71\pm 0.00$
BERTscore F1 $0.72\pm 0.01$ $0.71\pm 0.00$ $0.71\pm 0.01$ $0.72\pm 0.00$
BARTscore faithful $-2.75\pm 0.00$ $-2.7\pm 0.01$ $-2.71\pm 0.00$ $-2.78\pm 0.01$
BARTscore precision $-2.71\pm 0.01$ $-2.7\pm 0.00$ $-2.69\pm 0.00$ $-2.75\pm 0.00$
BARTscore recall $-2.77\pm 0.01$ $-2.79\pm 0.00$ $-2.8\pm 0.00$ $-2.76\pm 0.00$
BARTscore F1 $-2.71\pm 0.00$ $-2.72\pm 0.01$ $-2.72\pm 0.00$ $-2.73\pm 0.00$
guess native system prompt
overall guess guess correctly
native non-native native non-native
BLEU $0.42\pm 0.00$ $0.42\pm 0.00$ $0.45\pm 0.01$ $0.42\pm 0.00$
ROUGE-1 $0.37\pm 0.00$ $0.38\pm 0.01$ $0.39\pm 0.01$ $0.37\pm 0.01$
ROUGE-2 $0.13\pm 0.00$ $0.13\pm 0.00$ $0.13\pm 0.01$ $0.13\pm 0.01$
ROUGE-L $0.27\pm 0.00$ $0.27\pm 0.01$ $0.27\pm 0.01$ $0.27\pm 0.00$
BERTscore precision $0.72\pm 0.00$ $0.72\pm 0.00$ $0.71\pm 0.01$ $0.72\pm 0.01$
BERTscore recall $0.71\pm 0.00$ $0.71\pm 0.00$ $0.71\pm 0.01$ $0.71\pm 0.00$
BERTscore F1 $0.71\pm 0.00$ $0.72\pm 0.01$ $0.71\pm 0.01$ $0.71\pm 0.00$
BARTscore faithful $-2.78\pm 0.01$ $-2.78\pm 0.01$ $-2.75\pm 0.02$ $-2.74\pm 0.02$
BARTscore precision $-2.77\pm 0.00$ $-2.78\pm 0.00$ $-2.79\pm 0.06$ $-2.82\pm 0.02$
BARTscore recall $-2.76\pm 0.01$ $-2.74\pm 0.01$ $-2.76\pm 0.01$ $-2.76\pm 0.01$
BARTscore F1 $-2.74\pm 0.01$ $-2.74\pm 0.01$ $-2.76\pm 0.04$ $-2.76\pm 0.02$
Table 20: Overall average generative performance metrics and standard deviation over three runs for all generative tasks per group and prompt of GPT3.5.
(non-)native system prompt
correct information wrong information
native non-native native non-native
BLEU $0.47\pm 0.01$ $0.45\pm 0.01$ $0.45\pm 0.00$ $0.47\pm 0.00$
ROUGE-1 $0.40\pm 0.00$ $0.40\pm 0.00$ $0.39\pm 0.01$ $0.40\pm 0.00$
ROUGE-2 $0.15\pm 0.00$ $0.14\pm 0.01$ $0.14\pm 0.01$ $0.15\pm 0.00$
ROUGE-L $0.28\pm 0.01$ $0.28\pm 0.01$ $0.28\pm 0.00$ $0.29\pm 0.00$
BERTscore precision $0.72\pm 0.00$ $0.72\pm 0.01$ $0.72\pm 0.00$ $0.72\pm 0.00$
BERTscore recall $0.72\pm 0.00$ $0.71\pm 0.00$ $0.71\pm 0.00$ $0.72\pm 0.00$
BERTscore F1 $0.72\pm 0.00$ $0.72\pm 0.00$ $0.72\pm 0.00$ $0.72\pm 0.00$
BARTscore faithful $-2.95\pm 0.01$ $-2.88\pm 0.01$ $-2.87\pm 0.00$ $-2.96\pm 0.02$
BARTscore precision $-2.76\pm 0.01$ $-2.73\pm 0.01$ $-2.72\pm 0.01$ $-2.75\pm 0.01$
BARTscore recall $-2.79\pm 0.01$ $-2.8\pm 0.00$ $-2.81\pm 0.01$ $-2.77\pm 0.01$
BARTscore F1 $-2.75\pm 0.01$ $-2.74\pm 0.01$ $-2.74\pm 0.01$ $-2.74\pm 0.01$
guess native system prompt
overall guess guess correctly
native non-native native non-native
BLEU $0.46\pm 0.00$ $0.46\pm 0.00$ $0.48\pm 0.01$ $0.45\pm 0.01$
ROUGE-1 $0.40\pm 0.01$ $0.40\pm 0.00$ $0.39\pm 0.01$ $0.39\pm 0.02$
ROUGE-2 $0.15\pm 0.00$ $0.15\pm 0.00$ $0.13\pm 0.01$ $0.13\pm 0.02$
ROUGE-L $0.28\pm 0.01$ $0.29\pm 0.00$ $0.25\pm 0.01$ $0.27\pm 0.02$
BERTscore precision $0.72\pm 0.00$ $0.72\pm 0.00$ $0.69\pm 0.01$ $0.71\pm 0.01$
BERTscore recall $0.71\pm 0.00$ $0.72\pm 0.01$ $0.70\pm 0.01$ $0.71\pm 0.01$
BERTscore F1 $0.72\pm 0.00$ $0.72\pm 0.00$ $0.70\pm 0.01$ $0.71\pm 0.01$
BARTscore faithful $-2.92\pm 0.01$ $-2.91\pm 0.01$ $-2.58\pm 0.12$ $-2.76\pm 0.06$
BARTscore precision $-2.74\pm 0.01$ $-2.75\pm 0.01$ $-2.83\pm 0.04$ $-2.79\pm 0.07$
BARTscore recall $-2.80\pm 0.01$ $-2.79\pm 0.00$ $-2.74\pm 0.03$ $-2.84\pm 0.05$
BARTscore F1 $-2.75\pm 0.01$ $-2.74\pm 0.01$ $-2.77\pm 0.03$ $-2.79\pm 0.05$
Table 21: Overall average generative performance metrics and standard deviation over three runs for all generative tasks per group and prompt of GPT4o.
(non-)native system prompt
correct information wrong information
model native non-native native non-native
BLEU $0.45\pm 0.01$ $0.45\pm 0.01$ $0.45\pm 0.01$ $0.45\pm 0.01$
ROUGE-1 $0.41\pm 0.01$ $0.41\pm 0.00$ $0.41\pm 0.01$ $0.41\pm 0.00$
ROUGE-2 $0.16\pm 0.01$ $0.17\pm 0.00$ $0.16\pm 0.01$ $0.16\pm 0.01$
ROUGE-L $0.30\pm 0.01$ $0.30\pm 0.00$ $0.29\pm 0.01$ $0.30\pm 0.00$
BERTscore precision $0.72\pm 0.01$ $0.72\pm 0.00$ $0.72\pm 0.01$ $0.72\pm 0.00$
BERTscore recall $0.72\pm 0.01$ $0.72\pm 0.00$ $0.71\pm 0.00$ $0.72\pm 0.00$
BERTscore F1 $0.72\pm 0.01$ $0.72\pm 0.00$ $0.71\pm 0.01$ $0.72\pm 0.00$
BARTscore faithful $-2.89\pm 0.01$ $-2.88\pm 0.01$ $-2.87\pm 0.01$ $-2.91\pm 0.01$
BARTscore precision $-2.68\pm 0.01$ $-2.69\pm 0.01$ $-2.68\pm 0.01$ $-2.69\pm 0.01$
BARTscore recall $-2.70\pm 0.01$ $-2.68\pm 0.02$ $-2.71\pm 0.01$ $-2.68\pm 0.02$
BARTscore F1 $-2.67\pm 0.01$ $-2.66\pm 0.01$ $-2.67\pm 0.00$ $-2.66\pm 0.01$
guess native system prompt
overall guess guess correctly
metric native non-native native non-native
BLEU $0.44\pm 0.00$ $0.45\pm 0.00$ $0.34\pm 0.03$ $0.43\pm 0.00$
ROUGE-1 $0.39\pm 0.01$ $0.40\pm 0.01$ $0.38\pm 0.02$ $0.38\pm 0.00$
ROUGE-2 $0.14\pm 0.00$ $0.15\pm 0.00$ $0.17\pm 0.02$ $0.13\pm 0.00$
ROUGE-L $0.27\pm 0.01$ $0.28\pm 0.00$ $0.33\pm 0.02$ $0.27\pm 0.01$
BERTscore precision $0.71\pm 0.00$ $0.71\pm 0.00$ $0.74\pm 0.01$ $0.71\pm 0.00$
BERTscore recall $0.71\pm 0.00$ $0.71\pm 0.00$ $0.74\pm 0.01$ $0.71\pm 0.00$
BERTscore F1 $0.71\pm 0.00$ $0.71\pm 0.00$ $0.74\pm 0.01$ $0.71\pm 0.00$
BARTscore faithful $-2.97\pm 0.01$ $-2.97\pm 0.01$ $-3.33\pm 0.02$ $-2.92\pm 0.03$
BARTscore precision $-2.85\pm 0.01$ $-2.86\pm 0.01$ $-2.73\pm 0.02$ $-2.89\pm 0.03$
BARTscore recall $-2.76\pm 0.01$ $-2.73\pm 0.01$ $-2.63\pm 0.04$ $-2.77\pm 0.03$
BARTscore F1 $-2.77\pm 0.01$ $-2.77\pm 0.01$ $-2.64\pm 0.02$ $-2.80\pm 0.03$
Table 22: Overall average generative performance metrics and standard deviation over three runs for all generative tasks per group and prompt of Haiku.
(non-)native system prompt
correct information wrong information
model native non-native native non-native
BLEU $0.45\pm 0.00$ $0.45\pm 0.00$ $0.45\pm 0.00$ $0.45\pm 0.00$
ROUGE-1 $0.38\pm 0.00$ $0.38\pm 0.01$ $0.38\pm 0.00$ $0.39\pm 0.00$
ROUGE-2 $0.14\pm 0.00$ $0.14\pm 0.00$ $0.13\pm 0.00$ $0.14\pm 0.00$
ROUGE-L $0.27\pm 0.00$ $0.27\pm 0.00$ $0.27\pm 0.00$ $0.27\pm 0.01$
BERTscore precision $0.71\pm 0.00$ $0.71\pm 0.00$ $0.71\pm 0.00$ $0.71\pm 0.00$
BERTscore recall $0.71\pm 0.00$ $0.71\pm 0.00$ $0.71\pm 0.00$ $0.71\pm 0.00$
BERTscore F1 $0.71\pm 0.00$ $0.71\pm 0.00$ $0.71\pm 0.00$ $0.71\pm 0.00$
BARTscore faithful $-3.13\pm 0.00$ $-3.08\pm 0.01$ $-3.07\pm 0.01$ $-3.14\pm 0.01$
BARTscore precision $-2.87\pm 0.01$ $-2.86\pm 0.01$ $-2.86\pm 0.01$ $-2.87\pm 0.01$
BARTscore recall $-2.87\pm 0.00$ $-2.87\pm 0.01$ $-2.89\pm 0.00$ $-2.85\pm 0.01$
BARTscore F1 $-2.85\pm 0.00$ $-2.83\pm 0.01$ $-2.85\pm 0.00$ $-2.83\pm 0.01$
guess native system prompt
overall guess guess correctly
metric native non-native native non-native
BLEU $0.44\pm 0.01$ $0.45\pm 0.00$ $0.38\pm 0.02$ $0.46\pm 0.01$
ROUGE-1 $0.37\pm 0.01$ $0.38\pm 0.00$ $0.32\pm 0.02$ $0.39\pm 0.01$
ROUGE-2 $0.13\pm 0.00$ $0.13\pm 0.00$ $0.09\pm 0.02$ $0.13\pm 0.01$
ROUGE-L $0.26\pm 0.00$ $0.26\pm 0.00$ $0.23\pm 0.01$ $0.25\pm 0.01$
BERTscore precision $0.71\pm 0.00$ $0.71\pm 0.00$ $0.70\pm 0.02$ $0.70\pm 0.00$
BERTscore recall $0.71\pm 0.00$ $0.71\pm 0.00$ $0.70\pm 0.02$ $0.70\pm 0.00$
BERTscore F1 $0.71\pm 0.00$ $0.71\pm 0.00$ $0.70\pm 0.02$ $0.70\pm 0.00$
BARTscore faithful $-3.11\pm 0.01$ $-3.10\pm 0.00$ $-3.10\pm 0.08$ $-3.06\pm 0.01$
BARTscore precision $-2.88\pm 0.00$ $-2.89\pm 0.01$ $-2.96\pm 0.09$ $-2.93\pm 0.02$
BARTscore recall $-2.89\pm 0.00$ $-2.88\pm 0.01$ $-3.11\pm 0.1$ $-2.87\pm 0.01$
BARTscore F1 $-2.86\pm 0.00$ $-2.86\pm 0.01$ $-2.99\pm 0.1$ $-2.87\pm 0.01$
Table 23: Overall average generative performance metrics and standard deviation over three runs for all generative tasks per group and prompt of Sonnet.
(non-)native system prompt
correct information wrong information
model native non-native native non-native
BLEU $0.41\pm 0.00$ $0.42\pm 0.00$ $0.41\pm 0.01$ $0.42\pm 0.01$
ROUGE-1 $0.37\pm 0.01$ $0.38\pm 0.00$ $0.38\pm 0.01$ $0.38\pm 0.00$
ROUGE-2 $0.13\pm 0.01$ $0.14\pm 0.00$ $0.14\pm 0.01$ $0.14\pm 0.00$
ROUGE-L $0.26\pm 0.01$ $0.27\pm 0.00$ $0.27\pm 0.01$ $0.27\pm 0.00$
BERTscore precision $0.70\pm 0.00$ $0.70\pm 0.00$ $0.70\pm 0.00$ $0.70\pm 0.00$
BERTscore recall $0.69\pm 0.01$ $0.70\pm 0.00$ $0.69\pm 0.01$ $0.70\pm 0.00$
BERTscore F1 $0.69\pm 0.01$ $0.70\pm 0.00$ $0.70\pm 0.01$ $0.70\pm 0.00$
BARTscore faithful $-3.12\pm 0.01$ $-3.11\pm 0.01$ $-3.11\pm 0.01$ $-3.11\pm 0.01$
BARTscore precision $-2.93\pm 0.02$ $-2.92\pm 0.01$ $-2.91\pm 0.01$ $-2.93\pm 0.02$
BARTscore recall $-2.96\pm 0.02$ $-2.93\pm 0.02$ $-2.95\pm 0.02$ $-2.93\pm 0.02$
BARTscore F1 $-2.91\pm 0.02$ $-2.90\pm 0.02$ $-2.90\pm 0.02$ $-2.90\pm 0.02$
guess native system prompt
overall guess guess correctly
metric native non-native native non-native
BLEU $0.39\pm 0.01$ $0.39\pm 0.00$ $0.43\pm 0.02$ $0.32\pm 0.01$
ROUGE-1 $0.34\pm 0.00$ $0.34\pm 0.01$ $0.39\pm 0.01$ $0.30\pm 0.00$
ROUGE-2 $0.10\pm 0.00$ $0.10\pm 0.00$ $0.12\pm 0.02$ $0.07\pm 0.00$
ROUGE-L $0.23\pm 0.00$ $0.23\pm 0.00$ $0.24\pm 0.02$ $0.21\pm 0.00$
BERTscore precision $0.68\pm 0.00$ $0.68\pm 0.00$ $0.65\pm 0.01$ $0.65\pm 0.01$
BERTscore recall $0.68\pm 0.00$ $0.68\pm 0.00$ $0.67\pm 0.01$ $0.66\pm 0.00$
BERTscore F1 $0.68\pm 0.00$ $0.68\pm 0.00$ $0.66\pm 0.01$ $0.65\pm 0.01$
BARTscore faithful $-3.26\pm 0.01$ $-3.25\pm 0.01$ $-3.01\pm 0.03$ $-3.14\pm 0.07$
BARTscore precision $-3.16\pm 0.01$ $-3.17\pm 0.01$ $-3.14\pm 0.05$ $-3.36\pm 0.07$
BARTscore recall $-3.04\pm 0.01$ $-3.03\pm 0.00$ $-2.98\pm 0.04$ $-3.14\pm 0.03$
BARTscore F1 $-3.06\pm 0.01$ $-3.07\pm 0.01$ $-3.03\pm 0.04$ $-3.21\pm 0.04$
Table 24: Overall average generative performance metrics and standard deviation over three runs for all generative tasks per group and prompt of Qwen.
Appendix N Generative results generated label after annotation
Table 25 show the results for the overall average results for the generative tasks per model when letting the model first annotate the example and then gathering the output from this generated annotation as our ground truth label. As shown, this results in performance drops for both Qwen and GPT3.5. The larger models show similar performance to the experiments discussed in the paper.
GPT3.5 GPT4o Haiku Sonnet Qwen 7B
native non-native native non-native native non-native native non-native native non-native
BLEU 0.35 0.36 0.46 0.46 0.45 0.45 0.43 0.43 0.32 0.32
ROUGE-1 0.34 0.35 0.41 0.41 0.40 0.41 0.37 0.37 0.31 0.31
ROUGE-2 0.12 0.12 0.17 0.17 0.16 0.16 0.13 0.13 0.09 0.09
ROUGE-L 0.25 0.26 0.30 0.30 0.29 0.30 0.26 0.26 0.20 0.21
BERTscore precision 0.69 0.70 0.72 0.73 0.72 0.72 0.71 0.71 0.67 0.67
BERTscore recall 0.67 0.68 0.71 0.71 0.71 0.72 0.69 0.69 0.63 0.63
BERTscore F1 0.68 0.69 0.72 0.72 0.71 0.72 0.70 0.70 0.65 0.65
BARTscore faithful -2.80 -2.81 -2.95 -2.97 -2.88 -2.90 -3.12 -3.13 -3.13 -3.10
BARTscore precision -2.82 -2.81 -2.67 -2.68 -2.70 -2.71 -2.89 -2.90 -3.09 -3.07
BARTscore recall -2.98 -2.95 -2.74 -2.72 -2.68 -2.66 -2.98 -2.97 -3.37 -3.34
[…]

### v3

@@ Appendix I Results Sonnet different languages
Appendix K Classification results
Figure 11 shows the accuracy scores for the objective and subjective classification tasks per model when information about the nativeness of the prompt writer is added. We see how sonnet clearly performs differently than the other models.
Figure 11: Classification results per model and classification task when information about the nativeness of the prompt writer was added. We clearly see how Sonnet is highly influenced by this additional information.
Appendix L LLM as a judge: Generative results
Figure 12 shows how similar behavior is found across all three performance metrics per model. Moreover, Figure 13 shows the results per dataset for the generative results.
Figure 12: This Figure shows the performance of (western) native speakers and non-native speakers. We see how the highest performance for Coherence and is obtained for the western native group across all different models. The relevance scores show slightly less difference between groups, but the non-native and not western native group performs worse overall. The fluency scores are similar for all groups. We rescaled the results so that they range from 0 to 1.
Figure 13: This figure shows the overall performance across the three groups: (western) native speakers and non-native speakers. However, when looking into the coherence metric, we do see a preference for the western native group. The results show how there is no difference regarding fluency and only a slight performance difference when comparing the native categories with the non-native category for relevance.
Appendix M Additional Analysis
In this section, we include some extra analyses on the performance of the different groups within the non-native English speakers. More specifically, we add the results per level of English proficiency, as well as per frequency of English. We see that there are differences in performance across the different groups.
M.1 Classification results
For the classification results, we see a clear connection between performance and level of English, and frequency of usage of English. The groups with the highest levels of English also obtain better results. This is shown in Figures 14 and 15.
Figure 14: This figure shows the performance of English non-native speakers per self-reported level of English for the classification tasks.We adjusted the y-axis to range from 0.65 to 1 for clarity.
Figure 15: This figure shows the performance of English non-native speakers per self-reported frequency of English usage for the classification tasks. We adjusted the y-axis to range from 0.65 to 1 for clarity.
As we saw a performance difference, in terms of levels of English, we also compare the results when only taking into account level C1 and C2 non-native English speakers. The results are shown in Figure 16. Here, we still see the same order in performance as in Figure 3 was shown. However, now there is a clearer performance difference between the natives that are not western native and the non-native group.
Figure 16: This figure shows the performance of the three groups only including C2 and C1 level English speakers. We adjusted the y-axis to range from 0.65 to 1 for clarity.
M.2 Generative Results
For the generative tasks, however, we do not see clear differences in terms of frequency of English usage and performance, as shown in 17 and 18. Only the people with the lowest level of English proficiency perform better in terms of coherence, which is unexpected.
Figure 17: This figure shows the performance of English non-native speakers per self-reported level of English for the generative tasks. We rescaled the results so that they range from 0 to 1.
Figure 18: This figure shows the performance of English non-native speakers per self-reported frequency of English usage for the generative tasks. We rescaled the results so that they range from 0 to 1.
When analyzing the performance differences only for the groups with highest proficiency (C2 and C1), as shown in Figure 19, we see similar findings to Figure 4.
Figure 19: This figure shows the generative results only for the C2 and C1-level speakers per group.We rescaled the results so that they range from 0 to 1.
Appendix N Robustness analysis annotations
The dataset used in this paper is designed to be parallel, ensuring the same base samples for different annotator groups. We divided the dataset into multiple sets as shown in Table 5. As shown, the WN and NN groups have annotated all existing sets. The NWN group annotated most of the dataset, but did not annotate two out of the ten sets. Our dataset was constructed through random sampling and annotators from the three groups annotated the majority of sets. Below we provide a robustness check on the overlapping subsets.
Full dataset
Group
Objective
Classification
Subjective
Classification
WN 0.8661 0.7366
NWN 0.8487 0.7986
NN 0.8518 0.8039
Only Overlapping Sets
Group
Objective
Classification
Subjective
Classification
WN 0.8655 0.7366
NWN 0.8487 0.7986
NN 0.8492 0.8040
Table 13: Performance for the classification tasks on the full dataset and only considering overlapping samples.
[…]